for x being real when the expression (x-a)(x-b)/(x-c) assumes real values


ques of iit 1994


please give detailed solution

Grade:12th Pass

1 Answers

Chetan Mandayam Nayakar
12 years ago

(x-a)(x-b)/(x-c) =r, where r is real, (x-a)(x-b)=r(x-c),this is a quadratic equation in 'x', for x being real, the discriminant must be greater than or equal to zero